# CMB-HD as a Probe of Dark Matter on Sub-Galactic Scales

### v2

@@ III.1.1 Astrophysical Foregrounds
Figure 3: We show the residual extragalactic foreground power expected for CMB-HD at 90 GHz (top panel) and 150 GHz (bottom panel) after foreground cleaning (described in Section III.1.1), as well as the instrumental noise (gray dashed) and the lensed CMB temperature power spectrum (black). The residual power spectra of infrared (CIB) and radio sources are shown in green and purple, respectively, assuming that point sources detected at $5\sigma$ significance can be subtracted from the maps. In particular, we assume that radio sources with fluxes above 0.04 mJy can be detected at $5\sigma$ in the 90 GHz maps, corresponding to a flux limit of 0.03 mJy at 150 GHz. We assume that CIB sources with fluxes above 0.15 mJy can be detected at $5\sigma$ in the 280 GHz maps, corresponding to flux limits of 0.008 mJy and 0.03 mJy at 90 and 150 GHz, respectively. The residual CIB power also accounts for some source mis-subtraction due to uncertainty in the CIB spectral index. The residual thermal Sunyaev-Zel’dovich (tSZ) power (orange) is found by assuming that tSZ clusters detected above $3\sigma$ are subtracted from the maps. The CIB, radio, and tSZ power spectra were derived by [39]. The kinetic Sunyaev-Zel’dovich (kSZ) power, shown in blue, includes components from the epoch of reionization and from late times, calculated using templates from [40, 41] and [42], respectively; we do not assume that this signal will be removed. The total extragalactic foreground power is in red.
In this work, we account for the residual extragalactic foreground signal that will remain in the CMB-HD maps after foreground-cleaning procedures. Our foreground model includes contributions from the thermal and kinetic Sunyaev-Zel’dovich effects (tSZ and kSZ, respectively), the cosmic infrared background (CIB), and radio sources. The residual tSZ, CIB, and radio power spectra are obtained from [39], and we include the full power spectrum of the kinetic Sunyaev-Zel’dovich (kSZ) effect (with both reionization and late-time contributions) using a template from [42] for the late-time part, and from [40, 41] for the reionization part, normalized as described in Section III.6.222The model for the kSZ power spectrum used in this work differs from that in [39] and [38], by including both the late-time and reionization kSZ contributions, as opposed to just the latter. However, the contribution of small-scale $TT$ data to our lensing noise (Section III.1.3) and covariance matrix (Section III.1.4) is derived from the covariance matrix of [39], which only includes the reionization kSZ component. We show these residual foreground levels in Figure 3 for the 90 and 150 GHz channels, which are the frequencies that contain the most statistical weight in cosmological parameter analysis.
@@ III.1.2 Noise on CMB Power Spectra
We also include the effect of residual extragalactic foregrounds in the temperature data by adding the expected residual foreground power spectra for CMB-HD, described in Section III.1.1, to the temperature instrument noise spectra at 90 and 150 GHz. The coadded noise spectra, $N_{\ell}^{XY}$, for $XY\in\{TT,TE,EE,BB\}$ are obtained by calculating the beam-deconvolved instrumental noise power spectrum for each frequency channel (90 and 150 GHz), adding the residual foreground power to the beam-deconvolved instrumental noise for $TT$ for each frequency, and coadding the spectra of the two frequencies using inverse-noise weighting.
@@ III.1.3 Noise on CMB Lensing Spectra
The second estimator is the quadratic estimator of [52], which we refer to as the HDV estimator. It is designed to more optimally reconstruct small scales. Here, as in [38] and [39], we also use it to reduce bias to the lensing reconstruction on small-scales that arises due the presence of non-Gaussian foregrounds. This is done by filtering one CMB map to isolate large scales ($\ell<2000$), and the other CMB map to isolate only small scales ($\ell\in[5000,20000]$) (see [39] for details). We replace the noise from the H&O $TT$ estimator on scales $\ell>5000$ with the noise from the HDV $TT$ estimator, derived from the simulation-based covariance matrix of [39]. These simulations include extragalactic foregrounds that are correlated with the non-Gaussian CMB lensing map [39], using the the anticipated residual foreground levels for CMB-HD described in Section III.1.1 (with the exception of the kSZ effect, which only includes the reionization component as mentioned above).
@@ III.1.4 Construction of lensed and delensed covariance matrices
We follow a similar approach as [38] to calculate the covariance matrices for the lensed and delensed CMB spectra and CMB lensing spectrum in the multipole range $\ell\in[30,20000]$. Each of the lensed and delensed covariance matrices is a $5n_{\mathrm{bin}}\times 5n_{\mathrm{bin}}$ matrix, since five spectra are used in the forecasts. We denote each block by $\mathbb{C}_{\ell_{1}\ell_{2}}^{XY,WZ}$, where $XY$ and $WZ$ can each be one of $TT$, $TE$, $EE$, $BB$, and $\kappa\kappa$. We summarize the calculation below, which is discussed in more detail in [38], [51], [53], and [39].
The blocks of the covariance matrices are calculated analytically using the CLASS delens package [51] (unless otherwise specified). This calculation accounts for Gaussian and non-Gaussian terms (i.e., diagonal and off-diagonal elements, respectively), including:
  • The Gaussian variance between different CMB spectra, given by
  • The Gaussian variance of the CMB lensing power spectrum, given by
    where $XY$ and $WZ$ are each one of $TT$, $TE$, $EE$.
    where
The total covariance for the CMB $\times$ CMB blocks is then the sum of Eq. 1 and either 3 or 4, $\mathbb{C}_{\ell_{1}\ell_{2}}^{XY,WZ}=\mathbb{C}_{\ell_{1}\ell_{2},\mathrm{G}}^{XY,WZ}+\mathbb{C}_{\ell_{1}\ell_{2},\mathrm{NG}}^{XY,WZ}$. We assemble the full five-by-five block covariance matrix by joining the CMB×$\times$×CMB blocks with the CMB×$\times$×lensing (Eq. 6) and lensing×$\times$×lensing blocks (Eq. 2 plus the simulation-based covariance matrix from [39]).
[…]
@@ III.3 Calculating CMB and CMB Lensing Signal Spectra
While CAMB [57, 58] can calculate the lensed CMB and the CMB lensing power spectra in a $\Lambda$CDM model (with the option to include the effects of baryonic feedback using the HMcode2020 model [59]), it does not calculate these spectra for an arbitrary non-CDM model that changes the non-linear matter power spectrum. Thus, we calculate the CMB lensing power spectrum outside of CAMB as described below, and then input this into CAMB to get the lensed and delensed CMB power spectra. We show in Appendix A that for a $\Lambda$CDM model, our calculation agrees with the CAMB-only calculation to within 0.3% for $\ell\leq 20,000$, and to within 0.8% for $\ell\leq 40,000$.
@@ III.4.1 Lensing Signal-to-Noise Ratio per $k$ Mode
However, for visualization purposes, it is useful to see what comoving wavenumber $k$ in units of $h$ Mpc-1 the CMB-HD lensing signal is sensitive to (as we show in Figures 2 and 7). Below, we summarize how we approximate this, giving full details in Appendix B.
@@
We calculate the contribution to the lensing signal from a given redshift bin of width $\Delta z=0.5$, for redshifts $z\in[0,6]$. [^9] While the lensing signal is from all redshifts $z\in$ [0, 1100], we only go to $z=6$ in Figures 2 and 7 to keep our redshift bins small and ease the conversion from $\ell$ to $k$ at a given redshift. We note that the range $z\in[6,1100]$ contributes only 15% extra lensing signal. The $i^{\mathrm{th}}$ redshift bin ranges from $z_{i,\mathrm{min}}$ to $z_{i,\mathrm{max}}$, centered at $z_{i}$. We integrate Eq. 8 from $\chi(z_{i,\mathrm{min}})$ to $\chi(z_{i,\mathrm{max}})$ and then use Eq. 9 to calculate $C_{\ell}^{\kappa\kappa,z_{i}}$, which contains the lensing signal from only the $i^{\mathrm{th}}$ redshift bin. We use CAMB to obtain the lensed CMB spectra $C_{\ell}^{UV,z_{i}}$, for $UV\in\{TT,TE,EE,BB\}$, that were lensed by this $C_{\ell}^{\kappa\kappa,z_{i}}$. We then use the set of $C_{\ell}^{XY,z_{i}}$, for $XY\in\{TT,TE,EE,BB,\kappa\kappa\}$, in Eqs. 1 and 2 to calculate the diagonal elements of the covariance matrices for the five spectra, $\left(\sigma_{\ell}^{XY,z_{i}}\right)^{2}$. In Eqs. 1 and 2, $N_{\ell}^{XY}$ is the noise on the power spectrum, which we assume is the same for each redshift bin. The lensing SNR per multipole $\ell$ from the $i^{\mathrm{th}}$ redshift bin is then defined as
@@
While Eq. 16 underestimates the noise by neglecting the off-diagonal elements of the covariance matrix, it also underestimates the signal; when summing over redshift bins by adding in quadrature, the lensing cross-terms between redshift bins are neglected in the squared signal (see Appendix B). These two effects coincidentally nearly cancel, yielding a reasonable approximation to the SNR. For example, using Eq. 13 to calculate the lensing SNR from all the spectra yields 1947 (see Table 3 in Section IV), while using Eq. 14 and summing over all redshifts (to $z=1100$) and multipoles gives 1798, which is only slightly lower. [^10] When using only $z\in[0,6]$ as shown in Figure 7, Eq. 13 yields 1632, while Eq. 14 yields 1585 after summing over redshifts and multipoles, as shown in Table 6 in Appendix B.
@@ III.5 Including non-CDM Models
To include the effects of baryonic feedback on the non-linear matter distribution, we use the single-parameter baryonic feedback model of HMCode2020 [59] within CAMB to calculate the non-linear power spectrum, $P_{\Psi}^{\mathrm{CDM+baryons}}(k,z)$, directly (instead of applying a transfer function). This single-parameter model connects the various feedback parameters of hydrodynamic simulations to a single “AGN temperature” parameter, ${\rm{log_{10}}}(T_{\rm{AGN}}/{\rm{K}})$, which characterizes the overall strength of the feedback in the simulations. We then follow the procedure described in Section III.3 to calculate the power spectra with this model.
We show in Figure 5 the change in the $TT$ or lensing $\kappa\kappa$ power spectrum from a 1 keV WDM or a CDM+baryonic feedback model, relative to a CDM-only model. For the $TT$ spectrum, we only show the change due to the difference in the lensing effect; the change in the kSZ effect would show a similar behavior. We see both WDM and CDM+baryonic feedback models result in a suppression of power relative to a CDM-only model, however, the shape of the suppression differs considerably between the two models.
@@ III.7 Calculating the Fisher Matrix
For CMB-HD, the data vector $\hat{\vec{d}}$ is the set of binned CMB spectra (which could be lensed or delensed and includes the kSZ spectra) and the CMB lensing power spectrum. The covariance matrix is the binned covariance given by Eqs. 1 to 6, and the sum is taken over the bin centers. Each binned power spectrum is given by
@@
The Fisher matrix for the combination of CMB-HD and DESI is formed by taking the sum of their Fisher matrices, $F=F^{\mathrm{CMB}}+F^{\mathrm{BAO}}$. To calculate $F^{\mathrm{CMB}}$ for CMB-HD, we use Eq. 26 to calculate a Fisher matrix from the $TT,TE,EE,BB,\kappa\kappa$ power spectra in the range $\ell\in(30,20000)$ with the (binned) covariance matrix given by Eqs. 1 to 6. We then calculate a second Fisher matrix from the $TT$ power spectrum in the range $\ell\in(20000,40000)$, with a diagonal covariance matrix given by Eq. 1. We take the sum of these two Fisher matrices to obtain $F^{\mathrm{CMB}}$.
@@ IV.1 CMB-HD Lensing from $TT$, $TE$, $EE$, $BB$, and $\kappa\kappa$
In Figure 7, we illustrate which comoving scales and redshifts contribute to the CMB-HD lensing SNR given in the last row of Table 3. From Eq. 8, we see that lensing from a range of comoving wavenumbers, $k$, integrated over all redshifts, contribute to lensing at a given multipole, $\ell$. Thus we can only approximate the contribution to the lensing SNR for each redshift and wavenumber bin, as described in Section III.4.1 and Appendix B.
@@ IV.2 Parameter Forecasts
Figure 8: We show forecasts for a subset of the parameters varied in the Fisher analysis. The blue dotted contours and curves show the parameter constraints in a $\Lambda$CDM + $N_{\mathrm{eff}}$ + $\sum m_{\nu}$ + $A_{\mathrm{kSZ}}$ + $n_{\mathrm{kSZ}}$ model assuming CDM. The red dotted contours and curves show the forecasts for a 1 keV $\Lambda$WDM + $N_{\mathrm{eff}}$ + $\sum m_{\nu}$ + $m_{\mathrm{WDM}}$ + $A_{\mathrm{kSZ}}$ + $n_{\mathrm{kSZ}}$ model. We also show forecasts for these two models when including the effects of baryonic feedback (cyan contours and curves for CDM, pink contours and curves for WDM), with the strength of the feedback characterized by $\log_{10}(T_{\mathrm{AGN}}/\mathrm{K})$; we place a prior on this parameter from the anticipated cross-correlation of CMB-HD SZ and CMB lensing measurements as done in [38]. For the WDM + baryons model, we find no degeneracy between $\log_{10}(T_{\mathrm{AGN}}/\mathrm{K})$ and the WDM mass, $m_{\mathrm{WDM}}$, indicating that they can be separately distinguished.
Figure 9: We show the same forecasts as in Figure 8, now plotting the full set of varied parameters. For both CDM and WDM models, we marginalize over the baryonic feedback parameter, $\log_{10}(T_{\mathrm{AGN}}/\mathrm{K})$, with almost no increase in uncertainty on the other cosmological parameters (shown by the difference between dotted and solid contours of the same color). Table 4 lists the $1\sigma$ parameter uncertainties corresponding to this figure.

### v3

@@ III.1.1 Astrophysical Foregrounds
Figure 3: We show the residual extragalactic foreground power expected for CMB-HD at 90 GHz (top panel) and 150 GHz (bottom panel) after foreground cleaning (described in Section III.1.1 and Appendix A), as well as the instrumental noise (gray dashed) and the lensed CMB temperature power spectrum (black). The residual power spectra of infrared (CIB) and radio sources are shown in green and purple, respectively, assuming that point sources detected at $5\sigma$ significance can be subtracted from the maps. In particular, we assume that radio sources with fluxes above 0.04 mJy can be detected at $5\sigma$ in the 90 GHz maps, corresponding to a flux limit of 0.03 mJy at 150 GHz. We assume that CIB sources with fluxes above 0.15 mJy can be detected at $5\sigma$ in the 280 GHz maps, corresponding to flux limits of 0.008 mJy and 0.03 mJy at 90 and 150 GHz, respectively. The residual CIB power also accounts for some source mis-subtraction due to uncertainty in the CIB spectral index. The residual thermal Sunyaev-Zel’dovich (tSZ) power (orange) is found by assuming that tSZ clusters detected above $3\sigma$ are subtracted from the maps. The CIB, radio, and tSZ power spectra were derived by [39]. The kinetic Sunyaev-Zel’dovich (kSZ) power, shown in blue, includes components from the epoch of reionization and from late times, calculated using templates from [40, 41] and [42], respectively; we do not assume that this signal will be removed. The total extragalactic foreground power is in red.
In this work, we account for the residual extragalactic foreground signal that will remain in the CMB-HD maps after foreground-cleaning procedures. Our foreground model includes contributions from the thermal and kinetic Sunyaev-Zel’dovich effects (tSZ and kSZ, respectively), the cosmic infrared background (CIB), and radio sources. The residual tSZ, CIB, and radio power spectra are obtained from [39], and we include the full power spectrum of the kinetic Sunyaev-Zel’dovich (kSZ) effect (with both reionization and late-time contributions) using a template from [42] for the late-time part, and from [40, 41] for the reionization part, normalized as described in Section III.6.222The model for the kSZ power spectrum used in this work differs from that in [39] and [38], by including both the late-time and reionization kSZ contributions, as opposed to just the latter. However, the contribution of small-scale $TT$ data to our lensing noise (Section III.1.3) and covariance matrix (Section III.1.4) is derived from the covariance matrix of [39], which only includes the reionization kSZ component. We show these residual foreground levels in Figure 3 for the 90 and 150 GHz channels, which are the frequencies that contain the most statistical weight in cosmological parameter analysis. We describe our extragalactic foreground model in detail in Appendix A.
@@ III.1.2 Noise on CMB Power Spectra
We also include the effect of residual extragalactic foregrounds in the temperature data by adding the expected residual foreground power spectra for CMB-HD, described in Section III.1.1 and Appendix A, to the temperature instrument noise spectra at 90 and 150 GHz. The coadded noise spectra, $N_{\ell}^{XY}$, for $XY\in\{TT,TE,EE,BB\}$ are obtained by calculating the beam-deconvolved instrumental noise power spectrum for each frequency channel (90 and 150 GHz), adding the residual foreground power to the beam-deconvolved instrumental noise for $TT$ for each frequency, and coadding the spectra of the two frequencies using inverse-noise weighting.
@@ III.1.3 Noise on CMB Lensing Spectra
The second estimator is the quadratic estimator of [49], which we refer to as the HDV estimator. It is designed to more optimally reconstruct small scales. Here, as in [38] and [39], we also use it to reduce bias to the lensing reconstruction on small-scales that arises due the presence of non-Gaussian foregrounds. This is done by filtering one CMB map to isolate large scales ($\ell<2000$), and the other CMB map to isolate only small scales ($\ell\in[5000,20000]$) (see [39] for details). We replace the noise from the H&O $TT$ estimator on scales $\ell>5000$ with the noise from the HDV $TT$ estimator, derived from the simulation-based covariance matrix of [39]. These simulations include extragalactic foregrounds that are correlated with the non-Gaussian CMB lensing map [39], using the anticipated residual foreground levels for CMB-HD described in Section III.1.1 and Appendix A (with the exception of the kSZ effect, which only includes the reionization component as mentioned above).
@@ III.1.4 Construction of lensed and delensed covariance matrices
We follow a similar approach as [38] to calculate the covariance matrices for the lensed and delensed CMB spectra and CMB lensing spectrum in the multipole range $\ell\in[30,20000]$. Each of the lensed and delensed covariance matrices is a $5n_{\mathrm{bin}}\times 5n_{\mathrm{bin}}$ matrix, since five spectra are used in the forecasts. The full lensed or delensed covariance matrix includes diagonal terms to account for the Gaussian variance of the spectra, and off-diagonal terms to capture the non-Gaussian lensing-induced covariances between the spectra. We summarize the calculation in Appendix B, which is discussed in more detail in [38], [48], [50], and [39].
@@ III.3 Calculating CMB and CMB Lensing Signal Spectra
While CAMB [52, 53] can calculate the lensed CMB and the CMB lensing power spectra in a $\Lambda$CDM model (with the option to include the effects of baryonic feedback using the HMcode2020 model [54]), it does not calculate these spectra for an arbitrary non-CDM model that changes the non-linear matter power spectrum. Thus, we calculate the CMB lensing power spectrum outside of CAMB as described below, and then input this into CAMB to get the lensed and delensed CMB power spectra. We show in Appendix C that for a $\Lambda$CDM model, our calculation agrees with the CAMB-only calculation to within 0.3% for $\ell\leq 20,000$, and to within 0.8% for $\ell\leq 40,000$.
@@ III.4.1 Lensing Signal-to-Noise Ratio per $k$ Mode
However, for visualization purposes, it is useful to see what comoving wavenumber $k$ in units of $h$ Mpc-1 the CMB-HD lensing signal is sensitive to (as we show in Figures 2 and 7). Below, we summarize how we approximate this, giving full details in Appendix D.
@@
We calculate the contribution to the lensing signal from a given redshift bin of width $\Delta z=0.5$, for redshifts $z\in[0,6]$. [^6] While the lensing signal is from all redshifts $z\in$ [0, 1100], we only go to $z=6$ in Figures 2 and 7 to keep our redshift bins small and ease the conversion from $\ell$ to $k$ at a given redshift. We note that the range $z\in[6,1100]$ contributes only 15% extra lensing signal. The $i^{\mathrm{th}}$ redshift bin ranges from $z_{i,\mathrm{min}}$ to $z_{i,\mathrm{max}}$, centered at $z_{i}$. We integrate Eq. 2 from $\chi(z_{i,\mathrm{min}})$ to $\chi(z_{i,\mathrm{max}})$ and then use Eq. 3 to calculate $C_{\ell}^{\kappa\kappa,z_{i}}$, which contains the lensing signal from only the $i^{\mathrm{th}}$ redshift bin. We use CAMB to obtain the lensed CMB spectra $C_{\ell}^{UV,z_{i}}$, for $UV\in\{TT,TE,EE,BB\}$, that were lensed by this $C_{\ell}^{\kappa\kappa,z_{i}}$. We then use the set of $C_{\ell}^{XY,z_{i}}$, for $XY\in\{TT,TE,EE,BB,\kappa\kappa\}$, in Eqs. 23 and 24 to calculate the diagonal elements of the covariance matrices for the five spectra, $\left(\sigma_{\ell}^{XY,z_{i}}\right)^{2}$. In Eqs. 23 and 24, $N_{\ell}^{XY}$ is the noise on the power spectrum, which we assume is the same for each redshift bin. The lensing SNR per multipole $\ell$ from the $i^{\mathrm{th}}$ redshift bin is then defined as
@@
While Eq. 10 underestimates the noise by neglecting the off-diagonal elements of the covariance matrix, it also underestimates the signal; when summing over redshift bins by adding in quadrature, the lensing cross-terms between redshift bins are neglected in the squared signal (see Appendix D). These two effects coincidentally nearly cancel, yielding a reasonable approximation to the SNR. For example, using Eq. 7 to calculate the lensing SNR from all the spectra yields 1947 (see Table 3 in Section IV), while using Eq. 8 and summing over all redshifts (to $z=1100$) and multipoles gives 1798, which is only slightly lower. [^7] When using only $z\in[0,6]$ as shown in Figure 7, Eq. 7 yields 1632, while Eq. 8 yields 1585 after summing over redshifts and multipoles, as shown in Table 6 in Appendix D.
@@ III.5 Including non-CDM Models
To include the effects of baryonic feedback on the non-linear matter distribution, we use the single-parameter baryonic feedback model of HMCode2020 [54] within CAMB to calculate the non-linear power spectrum, $P_{\Psi}^{\mathrm{CDM+baryons}}(k,z)$, directly (instead of applying a transfer function). This single-parameter model connects the various feedback parameters of hydrodynamic simulations to a single “AGN temperature” parameter, ${\rm{log_{10}}}(T_{\rm{AGN}}/{\rm{K}})$, which characterizes the overall strength of the feedback in the simulations. While in general baryonic physics models can be quite complicated, a number of publications have suggested that the main effects can be characterized by a few parameters, or even as suggested by [54], by one effective parameter via which several others can be characterized (see, e.g. Table 5 in [54]). We then follow the procedure described in Section III.3 to calculate the power spectra with this single-parameter model.
We show in Figure 5 the change in the $TT$ or lensing $\kappa\kappa$ power spectrum from a 1 keV WDM or a CDM+baryonic feedback model, relative to a CDM-only model. For the $TT$ spectrum, we only show the change due to the difference in the lensing effect; the change in the kSZ effect would show a similar behavior. We see both WDM and CDM+baryonic feedback models result in a suppression of power relative to a CDM-only model, however, the shape of the suppression differs considerably between the two models. A wide range of baryonic feedback models and simulations that extend beyond $k\sim 10$ $h$/Mpc, such as those considered in [54, 63, 64, 65, 66], seem to share a similar suppression at moderate scales followed by an enhancement at smaller scales (“U-shape”). A WDM model, on the other hand, follows a scale-dependent suppression that increases steadily towards smaller scales, which is not easily mimicked by baryonic feedback. Therefore, while the details of each kind of effect can be more complicated, Figure 5 is still likely representative of CMB-HD’s ability to distinguish between the effects of baryonic feedback and alternative dark matter.
@@ III.7 Calculating the Fisher Matrix
For CMB-HD, the data vector $\hat{\vec{d}}$ is the set of binned CMB spectra (which could be lensed or delensed and includes the kSZ spectra) and the CMB lensing power spectrum. The covariance matrix is the binned covariance given by Eqs. 23 to 28, and the sum is taken over the bin centers. Each binned power spectrum is given by
@@
The Fisher matrix for the combination of CMB-HD and DESI is formed by taking the sum of their Fisher matrices, $F=F^{\mathrm{CMB}}+F^{\mathrm{BAO}}$. To calculate $F^{\mathrm{CMB}}$ for CMB-HD, we use Eq. 20 to calculate a Fisher matrix from the $TT,TE,EE,BB,\kappa\kappa$ power spectra in the range $\ell\in(30,20000)$ with the (binned) covariance matrix given by Eqs. 23 to 28. We then calculate a second Fisher matrix from the $TT$ power spectrum in the range $\ell\in(20000,40000)$, with a diagonal covariance matrix given by Eq. 23. We take the sum of these two Fisher matrices to obtain $F^{\mathrm{CMB}}$.
@@ IV.1 CMB-HD Lensing from $TT$, $TE$, $EE$, $BB$, and $\kappa\kappa$
In Figure 7, we illustrate which comoving scales and redshifts contribute to the CMB-HD lensing SNR given in the last row of Table 3. From Eq. 2, we see that lensing from a range of comoving wavenumbers, $k$, integrated over all redshifts, contribute to lensing at a given multipole, $\ell$. Thus we can only approximate the contribution to the lensing SNR for each redshift and wavenumber bin, as described in Section III.4.1 and Appendix D.
@@ IV.2 Parameter Forecasts
Figure 8: We show our forecasts including the full set of varied parameters. The blue dotted contours and curves show the parameter constraints in a $\Lambda$CDM + $N_{\mathrm{eff}}$ + $\sum m_{\nu}$ + $A_{\mathrm{kSZ}}$ + $n_{\mathrm{kSZ}}$ model assuming CDM. The red dotted contours and curves show the forecasts for a 1 keV $\Lambda$WDM + $N_{\mathrm{eff}}$ + $\sum m_{\nu}$ + $m_{\mathrm{WDM}}$ + $A_{\mathrm{kSZ}}$ + $n_{\mathrm{kSZ}}$ model. We also show forecasts for these two models when including the effects of baryonic feedback (cyan contours and curves for CDM, pink contours and curves for WDM), with the strength of the feedback characterized by $\log_{10}(T_{\mathrm{AGN}}/\mathrm{K})$; we place a prior on this parameter from the anticipated cross-correlation of CMB-HD SZ and CMB lensing measurements as done in [38]. Table 4 lists the $1\sigma$ parameter uncertainties corresponding to this figure.
Figure 9: We show forecasts for a subset of the parameters varied in the Fisher analysis. For both CDM and WDM models, we marginalize over the baryonic feedback parameter, $\log_{10}(T_{\mathrm{AGN}}/\mathrm{K})$, with almost no increase in uncertainty on the other cosmological parameters (shown by the difference between dotted and solid contours of the same color). For the WDM + baryons model, we find no degeneracy between $\log_{10}(T_{\mathrm{AGN}}/\mathrm{K})$ and the WDM mass, $m_{\mathrm{WDM}}$, indicating that they can be separately distinguished.